Tyrotek Tyrotek Analog Computer
- Unit Log -

Model:not specified
Year of Manufacture:1968
Year of Manufacture Source:"Aug 68" on plastic label on front, date stamps of 6713 to 6734 on power supply filter capacitors.
Source:UBC SERF (Surplus Equipment Recycling Fund)
Date of Receipt:2000~
Other Identification:"Non Linear Mtg." in foil on some of the printed circuit boards.

2009 Jan: Reverse engineered to produce schematic.
2009 Jan: - Fully dismantled for cleaning.
- White powder had grown on various screws, nuts, washers, spacers, etc.
- Wires between the main and meter compartments had previously been damaged or cut and spliced where they go through the divider.
2009 Jan: Original power connector replaced with IEC connector.
2009 Jan 24: - Reassembly complete.
- Hardware replaced with primarily stainless steel items.
- Damaged wires replaced or fixed and some rewiring done around the meter compartment.
2009 Jan 24: Powered up with one ampliier and one multiplier installed. Pots and null divider work, simple summation with amp 02 functional.
2009 Jan 29: Remaining modules installed. Test program wired up to test all amps and overflow lamps. Lamp for A17 does not work, swapped with A20 which is unused. Some inconsistencies in operation of other lamps (A1,3,4,5,7), but meter indicates output from amps to be OK.
2009 Feb 02: Plug-wires cleaned. Light oil drop placed in ball-bearing of some units which were binding, those lubricated marked with a black dot on the orange hood. Black, yellow, red, blue wires with plastic insulation came with analog computer. Grey wires were from another source and are bi-plugs cut apart.
2011 Jan: Communication with Jim Pepper and family.
2011 Jan 12: Attempt to understand multiplier operation based on Jim's hint of time division operation. Multiplier examined with scope. Q5 & Q6 test points on MD adjunct show oscillation at 120 KHz. +-Ref => pot => amp => Y input produces variable duty cycle at Q5, Q6 test points.
2011 Jan 13: Other multiplier tested, not functional. On multiplier sub-board, Q2 Drain shows no signal. Q1 and Q2 removed for testing, Q1 appears OK with simple switching action test, Q2 shows no action. Replacement needed for Q2, P-channel enhancement MOSFET.

Tyrotek Analog Computer
2009 Jan